f(x)=7x+2 Find f(x) when x=0andwhen,x=1
Answer by jim_thompson5910(35256)

f(x)=7x+2

f(0)=7(0)+2

f(0) = 0+2

f(0) = 2

-------------------------------------------------------

f(x)=7x+2

f(0)=7(1)+2

f(0) = 7+2

f(0) = 9


If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim